OpenStudy (anonymous):

Find the points of intersection of y= x^2 - 3x and y= -2x^2 -5x-24

OpenStudy (zarkon):

there aren't any real intersection points

OpenStudy (anonymous):

how comes?

OpenStudy (anonymous):

Set them equal to each other so -2x^2-5x-24=x^2-3x 3x^2+2x+24=0. However, this is not factorable so there are no real points of intersection.

OpenStudy (zarkon):

try graphing them... you will see that they never intersect

OpenStudy (anonymous):

it does intersect for the first one it may be (4,4) and the second one (-3, 18) ??

OpenStudy (zarkon):

no...they don't intersect. if you think they intersect them maybe you wrote the problem incorrectly.

OpenStudy (anonymous):

oh yahh they dont intersect! haha thank you!